Circular No. 7108

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999aa IN NGC 2595 Ron Arbour, South Wonston, Hants, U.K., reports his discovery of an

apparent supernova (mag ∼ 15.5) in the galaxy NGC 2595 on an unfiltered CCD image (limiting mag ∼ 17.5) taken on Feb. 11.0166 UT using a 0.3m f/6.3 reflector in the course of his supernova patrol. M. Armstrong confirmed the new object’s existence via two exposures (the first at Feb. 11.0461 UT), measuring its position as α = 8^{h}27^{m}42^{s}.03, δ = +21^{o}29′14′′.8 (equinox 2000.0), which Arbour notes to be 1′′ east and 28′′ north of the center of NGC 2595. Nothing is present at this position on six images taken by Arbour in 1998 Feb. and Mar. (limiting mag ∼ 17.5), nor is it present in the Digitized Sky Survey. Arbour notes that a CCD image obtained by M. Schwartz with the Tenagra 0.35-m patrol telescope on Feb. 12.07 yields an offset of 2′′ east, 27′′ north, while an image by Schwartz obtained on 1998 Nov. 26.45 shows no object at this location.

A. V. Filippenko, W. D. Li, and D. C. Leonard, University of California at Berkeley, write: “CCD spectra (range 330–1000 nm) obtained on Feb. 12 UT with the Lick Observatory 3-m Shane reflector show that SN 1999aa is a peculiar type-Ia supernova, very similar to SN 1991T (Filippenko et al. 1991, Ap.J. 384, L15). The Si II absorption at 620 nm is very weak, but Fe III lines near 430 and 500 nm are prominent; unlike the case of SN 1991T, however, the absorption trough centered at 375 nm (due to Ca II?) is strong. Overall, the spectrum suggests an age of ∼ 6 days before maximum brightness. SN 1999aa is not visible in an unfiltered image of NGC 2595 obtained on Feb. 5.3 (limiting mag ∼ 19) with the 0.8-m Katzman Automatic Imaging Telescope in the course of the Lick Observatory Supernova Search (cf. IAUC 6627). The recession velocity of the host galaxy, determined from narrow Hα and [N II] emission, is 4400 km/s.”

RINGS OF NEPTUNE C. Roddier, F. Roddier, J. E. Graves, O. Guyon, and M. J. Northcott,

Institute for Astronomy, University of Hawaii (UH); and B. Sicardy, Paris Observatory, report: “Infrared (1.72-μm) images of Neptune taken on 1998 July 6 with the UH adaptive optics system (Hokupa’a) mounted on the Canada-France-Hawaii 3.6-m telescope have been studied further. In addition to the observations of Neptune arcs reported on IAUC 7051, evidence is now found for another ring structure even closer to Neptune. Its radial distance is that of the Le Verrier ring. It is visible only on the west ansa, and its brightness is roughly twice that of the Adams arcs.”

1999 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green